OpenStudy (anonymous):

The Hudson River flows at a rate of 3 miles per hour. A patrol boat travels 60 miles upriver, and returns in a total time of 9 hours. What is the speed of the boat in still water?

OpenStudy (anonymous):

(Distance)/(Time) = Rate 1. 60/(R-3) + 60/(R+3) = 9 solve for R 9R^2 - 120R - 81 = 0 R = 13.98 mph

OpenStudy (anonymous):

thank you